# Sum frequency generation from real-time simulations in two-dimensional crystals

### v1

@@ 5.1 SFG and DFG in monolayer h-BN
In Fig. 4 we report the SFG and DFG spectra for h-BN both at the IPA and TD-aGW approximation level. For SFG, the line $\omega_{1}=\omega_{2}$ corresponds to the SHG, already calculated in Ref. [13].In Fig. 4, we report the SFG and DFG spectra for h-BN both at the IPA and TD-aGW level. At the IPA level, the SFG spectrum (panel (a)) of h-BN is dominated by two-photon resonances with single-particle transitions between 4-6 eV (negative slopes bands). In particular, the lowest-energy band corresponds to the transition from the valence band minimum to the conduction band maximum. The one-photon transitions (vertical and horizontal bands) are much weaker though a significant enhancement is observed for part of the spectra both resonant with one- and two-photon. In particular, this portion of the SFG spectrum is twice as intense as the SHG (the $\omega_{1}=\omega_{2}$ diagonal). The DFG spectrum (panel (b)) is dominated by the resonant optical rectification between 4-6 eV ($\omega_{1}=\omega_{2}$). Further, one-photon resonances (vertical and horizontal bands) are visible, again enhanced when two-photon resonances are also present. These results can be straightforwardly related to the IPA absorption spectrum (see e.g. [13]), which presents a broad peak between 4-6 eV).
@@
As expected, the addition of the electron-hole interaction drastically changes the SFG/DFG spectra [^2] Note that the shift of the onset is due to the addition of the scissor operator in Table 1 partially compensated by the exciton binding energy.. Sharper and much stronger (note the different color scale) features appear in the TD-aGW spectra, corresponding to the $E_{1},E_{2}$ excitonic peaks at around 6.1 eV and 7 eV.[32]. Two-photon resonances with the two excitons are clearly visible in the SFG spectrum (negative slope lines in panel (c)) and DFG spectrum (positive slope lines starting at the exciton energy in panel (d)). One-photon resonances are also visible (as vertical/horizontal lines). The responses are significantly enhanced in the SFG (DFG) when one laser is resonant with $E_{1}$ ($E_{2}$) and the second laser with $E_{2}-E_{1}$. These spectral features correspond to the first and third terms in Eq. (6) and provide a measure of the strength of exciton-exciton transitions. In the DFG spectrum, strong features are visible as well on the diagonal, corresponding to the exciton-resonant optical rectification and when one laser is resonant with one exciton and the second with the other (corresponding to the second and fourth term in Eq. (6)).
@@
Figure 5: SFG/DFG spectra for MoS2 in panels (a) and (b) at the independent particle level and in panel (c), (d) at the TD-aGW level. These heatmaps have been generated using a frequency grid of $\omega_{1}\times\omega_{2}=96\times 96$ points for the IPA and $\omega_{1}\times\omega_{2}=72\times 72$ for the TD-aGW. For each frequency pair a real-time simulation was run and the output signal processed. .
5.2 SFG and DFG in monolayer MoS2
@@
In Fig. 5, we report the SFG and DFG spectra for the monolayer MoS2 at the IPA and TD-aGW level. When compared with h-BN, the differences between the IPA and TD-aGW spectra are less striking, as already observed for the absorption [33] and the SHG [13]. Similarly to what observed for the SHG[34, 13, 35], a significant enhancement is seen at resonances with the C exciton ($\approx 3$ eV), while the weaker the A and B excitons ($\approx 2.2$ eV, as spin-orbit coupling is not included the peaks are degenerate) show minimal excitonic enhancement. The SFG spectrum shows a strong two-photon resonance with the C exciton (negative slope line) while the DFG spectrum shows a strong one-photon resonance (vertical/horizontal line) and a strong exciton resonant optical rectification. Though less evident than for h-BN, there is an enhancement in the intensity in correspondence of exciton-exciton transitions. In the SFG (panel (c)), the signal is enhanced when one laser is resonant with the A/B exciton (about 2.2 eV) while the frequency of the second matches the energy difference between the C and A/B excitons. In the DFG (panel (d)), the signal is enhanced when one laser is resonant with the C exciton (about 3 eV) while the frequency of the second corresponds to the energy difference between the C and A/B excitons. Simulations results align with existing measurements. SFG has been observed [36] at 2.9 eV when laser fields at 1.2 eV and 1.9 eV were injected. The DFG has been observed [7] fixing one laser at 3.06 eV and varying the second between 0.79 and 0.95 eV showing an enhancement of the signal between 2.1-2.2 eV that the authors attributed to excitonic effects in this region.
@@
Figure 6: Calculated THz field-induced second-harmonic generation of bilayer h-BN with $\nu$=10 THz at the IPA level. Each curve consists of 192 frequency steps between $\omega=1-6~{}\mathrm{eV}$.
@@ 5.3 Field-Induced Second-Harmonic Generation
In Fig. 6 we report the third order susceptibility corresponding to the FI-SHG in bilayer h-BN. For each frequency $\omega$, we run a real-time simulation in presence of the THz pump ($\nu$ = 10 THz). The response functions reported in the figure correspond to two possible experimental configurations, a pump and probe in $y$ direction ($\chi^{(3)}_{2222}$), and pump in $x$ and probe in $y$ direction ($\chi^{(3)}_{1221}$). In both configurations, the susceptibility shows a strong resonance at half of the gap, around 2-3 eV, similar to the standard SHG in monolayer h-BN [13]. We found that the intensity of the response when pump and probe are parallel is higher than the perpendicular configuration. The order of magnitude of the $\chi^{(3)}$ is comparable with that of bulk ferroelectric oxides which are known for their excellent nonlinear properties [37]. This result implies that two-dimensional crystals can be used as a detector for THz radiation [38].
@@ 6 Conclusions
In this work, we present a computational framework to study sum/difference frequency generation by means of real-time simulations in presence of multiple laser fields. With multiple fields, the challenge is the signal processing required to extract the nonlinear susceptibilities. In particular, using a discrete Fourier transform approach may require very long and thus computationally costly simulations. We found that approaches based either on the singular value decomposition and the least squares optimization give accurate results with short sampling time and allow to reduce significantly the simulation time. These approaches enable the calculation of second-order response functions, as SFG/DFG, and higher nonlinear response functions, as FI-SHG, including excitonic effects within many-body theory. For the system studied, monolayer h-BN and MoS2, we showed that including excitonic effects in SFG/DFG spectra is critical. In both materials, we predict strong features corresponding to exciton transitions, as recently experimentally observed for another layered material [8]. Further, the results on FI-SHG for bilayer demonstrate that the approach can be used to predict and interpret nonlinear terahertz spectroscopy of solids [39]. Finally, the approach presented can be coupled to atomic vibrations by using finite displacement methods[40, 41] opening the way to simulate other spectroscopic techniques such as Coherent Anti-Stokes Raman spectroscopy (CARS). The latter is a powerful nonlinear optical technique for probing vibrational modes in molecular and solid-state structures. CARS involves two laser beams exciting a vibrational state, and a third beam generating a coherent Anti-Stokes signal, allowing for high resolution imaging [42]. CARS can be seen as a combination of SFG and DFG processes, and therefore the method shown in this study could be employed to study the nonresonant CARS response. Since the pure nonresonant CARS is not directly available experimentally [43], the ab initio real-time simulation is a promising feature to support CARS measurements.
@@ Acknowledgments
The authors gratefully acknowledge financial support from the Deutsche Forschungsgemeinschaft (DFG) through the research group FOR5044 (ID: 426703838; http://www.For5044.de). Calculations for this research were conducted on the Lichtenberg high-performance computer of the TU Darmstadt and at the Höchstleistungsrechenzentrum Stuttgart (HLRS). The authors furthermore acknowledge the computational resources provided by the HPC Core Facility and the HRZ of the Justus-Liebig-Universität Gießen. C.A. acknowledges B. Demoulin and A. Saul for the management of the computer cluster Rosa. C.A. acknowledges the COLIBRI No. ANR-22-CE30-0027 ANR project. C.A. and M.G. acknowledge funding from European Research Council MSCA-ITN TIMES under grant agreement 101118915 and from AMUtech. M.G. acknowledges funding from the UKRI Horizon Europe Guarantee funding scheme (EP/Y032659/1).

### v2

@@ 5.1 SFG and DFG in h-BN monolayer
In Fig. 4 we report the SFG and DFG spectra for h-BN both at the IPA and TD-aGW approximation level. For SFG, the line $\omega_{1}=\omega_{2}$ corresponds to the SHG, already calculated in Ref. [14]. At the IPA level, the SFG spectrum (Fig. 4(a)) of h-BN is dominated by two-photon resonances with single-particle transitions between 4-6 eV (negative slope bands). In particular, the lowest-energy band corresponds to the transition from the valence band minimum to the conduction band maximum. The one-photon transitions (vertical and horizontal bands) are much weaker though a significant enhancement is observed for part of the spectra both resonant with one- and two-photon absorption. In particular, this portion of the SFG spectrum is twice as intense as the SHG (the $\omega_{1}=\omega_{2}$ diagonal). The DFG spectrum (Fig. 4(b)) is dominated by the resonant optical rectification between 4-6 eV ($\omega_{1}=\omega_{2}$). Further, one-photon resonances (vertical and horizontal bands) are visible, again enhanced when two-photon resonances are also present. These results can be straightforwardly related to the IPA absorption spectrum (see e.g. Ref. [14], which presents a broad peak between 4-6 eV). This broad peak can be found in the calculated absorption spectrum in Fig. 5(a).
Figure 5: Calculated imaginary dielectric function of (a) h-BN and (b) at the IPA (blue dashed line) and TD-aGW (orange solid line) level. The excitonic peaks $E_{1}$ and $E_{2}$ of h-BN are located at around 6.1 eV and 7 eV, respectively (panel (a)). The degenerate excitonic peak A/B as well as C can be seen at around 2.2 eV and 3 eV, respectively (panel (b)).
@@
As expected, the addition of the electron-hole interaction drastically changes the SFG/DFG spectra [^4] Note that the shift of the onset is due to the addition of the scissor operator in Table 1 partially compensated by the exciton binding energy.. Sharper and much stronger (note the different color scale) features appear in the TD-aGW spectra, corresponding to the $E_{1},E_{2}$ excitonic peaks at around 6.1 eV and 7 eV [36]. These two excitonic peaks can also be identified in the calculated absorption spectrum at TD-aGW level in Fig. 5(a). Two-photon resonances with the two excitons are clearly visible in the SFG spectrum (negative slope lines in Fig. 4(c)) and DFG spectrum (positive slope lines starting at the exciton energy in Fig. 4(d)). One-photon resonances are also visible (as vertical/horizontal lines). The responses are significantly enhanced in the SFG (DFG) when one laser is resonant with $E_{1}$ ($E_{2}$) and the second laser with $E_{2}-E_{1}$. These spectral features correspond to the first and third terms in Eq. (8) and provide a measure of the strength of exciton-exciton transitions. In the DFG spectrum, strong features are visible as well on the diagonal, corresponding to the exciton-resonant optical rectification, as well as when one laser is resonant with one exciton and the second with the other (corresponding to the second and fourth term in Eq. (8)).
The results presented here on the h-BN monolayer are a proof of concept of our methodology, showing that resonances with strongly bound excitons are important in both SFG/DFG spectra. Due to the large band gap of h-BN, the region of interest for SFG/DFG is difficult to sample and to our knowledge there are no experimental measurements.
@@
Figure 6: SFG/DFG spectra for MoS2 in panels (a) and (b) at the independent particle level and in panel (c), (d) at the TD-aGW level. These heatmaps have been generated using a frequency grid of $\omega_{1}\times\omega_{2}=96\times 96$ points for the IPA and $\omega_{1}\times\omega_{2}=72\times 72$ for the TD-aGW. For each frequency pair a real-time simulation was run and the output signal processed.
5.2 SFG and DFG in MoS2 monolayer
@@
In Fig. 6, we report the SFG and DFG spectra for the MoS2 monolayer at the IPA and TD-aGW level. When compared with h-BN, the differences between the IPA and TD-aGW spectra are less striking, as already observed for the absorption [37] and the SHG [14]. Similarly to what was observed for the SHG[38, 14, 39], a significant enhancement is seen at resonances with the C exciton ($\approx 3$ eV), while the weaker A and B excitons ($\approx 2.2$ eV, as spin-orbit coupling is not included the peaks are degenerate) show minimal excitonic enhancement. These excitonic peaks are also visible in the linear absorption spectrum as shown in Fig. 5(b). The SFG spectrum shows a strong two-photon resonance with the C exciton (negative slope line) while the DFG spectrum shows a strong one-photon resonance (vertical/horizontal line) and a strong exciton resonant optical rectification. Though less evident than for h-BN, there is an enhancement in the intensity in correspondence of exciton-exciton transitions. In the SFG (Fig. 6(c)), the signal is enhanced when one laser is resonant with the A/B exciton (about 2.2 eV) while the frequency of the second matches the energy difference between the C and A/B excitons. In the DFG (Fig. 6(d)), the signal is enhanced when one laser is resonant with the C exciton (about 3 eV) while the frequency of the second corresponds to the energy difference between the C and A/B excitons.
SFG and DFG were measured in mono- [40, 7, 41, 5] and few-layers [4] MoS2. In the presence of metal substrates [40], excitonic resonances were shown to be strongly attenuated, and their position shifted due to gap renormalization. These effects are beyond the methodology presented in this manuscript. Other measurements are performed in a pump-probe configuration with a delay between the pump and probe [40], which in our case requires the inclusion of dephasing effects that are beyond the scope of the present work. Finally, for insulating substrates and synchronized pump and probe, our simulation results are in agreement with existing measurements. In particular, SFG has been observed [41] at 2.9 eV when laser fields at 1.2 eV and 1.9 eV were injected. The DFG was observed [7] by fixing one laser at 3.06 eV and varying the second between 0.79 and 0.95 eV, showing an enhancement of the signal between 2.1-2.2 eV, which the authors attributed to excitonic effects in this region. Our results support this interpretation.
@@
Figure 7: Calculated THz field-induced second-harmonic generation of the h-BN bilayer with $\nu$=10 THz at the IPA level. Each curve consists of 192 frequency steps between $\omega=1-6~\mathrm{eV}$.
@@ 5.3 Field-induced second-harmonic generation
[…]
@@ 6 Conclusions
In this work, we present a computational framework to study sum/difference frequency generation by means of real-time simulations in the presence of multiple laser fields [^5] After submission, we became aware of a newly published study proposing an alternative approach to the one we have developed [44].. With multiple fields, the challenge is the signal processing required to extract the nonlinear susceptibilities. In particular, using a discrete Fourier transform approach may require very long and thus computationally costly simulations. We found that approaches based either on the singular value decomposition and the least squares optimization give accurate results with short sampling time and allow to reduce significantly the simulation time. These approaches enable the calculation of second-order response functions, such as SFG/DFG, and higher nonlinear response functions, as FI-SHG, including excitonic effects within many-body theory. For the studied systems, h-BN monolayer and MoS2, we showed that including excitonic effects in SFG/DFG spectra is critical. In both materials, we predict strong features corresponding to exciton transitions, as recently experimentally observed for another layered material [9]. Further, the results on FI-SHG for the h-BN bilayer demonstrate that the approach can be used to predict and interpret nonlinear terahertz spectroscopy of solids [45]. Finally, the presented approach can be coupled to atomic vibrations using finite displacement methods [46, 47] opening the way to simulate other spectroscopic techniques such as coherent anti-Stokes Raman spectroscopy (CARS). The latter is a powerful nonlinear optical technique for probing vibrational modes in molecular and solid-state structures. CARS involves two laser beams exciting a vibrational state, and a third beam generating a coherent anti-Stokes signal, allowing for high resolution imaging [48]. CARS can be seen as a combination of SFG and DFG processes, and therefore the method shown in this study could be employed to study the nonresonant CARS response. Since the pure nonresonant CARS is not directly available experimentally [49], the ab initio real-time simulation is a promising feature to support CARS measurements.
@@ Acknowledgments
The authors gratefully acknowledge financial support from the Deutsche Forschungsgemeinschaft through the FOR5044 research group (ID: 426703838; http://www.For5044.de). Calculations for this research were conducted on the Lichtenberg high-performance computer of the TU Darmstadt and at the Höchstleistungsrechenzentrum Stuttgart (HLRS). The authors furthermore acknowledge the computational resources provided by the HPC Core Facility and the HRZ of the Justus-Liebig-Universität Gießen. C.A. acknowledges B. Demoulin and A. Saul for the management of the computer cluster Rosa. C.A. acknowledges the French “Agence Nationale de la Recherche (ANR)” for financial support (Grant Agreement No. ANR-22-CE30-0027). C.A. and M.G. acknowledge funding from European Research Council MSCA-ITN TIMES under grant agreement 101118915 and from AMUtech. M.G. acknowledges funding from the UKRI Horizon Europe Guarantee funding scheme (EP/Y032659/1).